2b) A plane wave is described by an electric field

180πcos10 (2π 100 ).

8 E  t  y

What is its (i) frequency in Hz, (ii) wave number, (iii) phase velocity,

(iv) wavelength, and (v) direction of propagation?


1
Expert's answer
2021-03-29T09:06:30-0400
"f=100\ Hz\\k=8\ m^{-1}\\v=\frac{2\pi(100)}{8}=78.5\frac{m}{s}\\\lambda=0.125\ m\\"

Direction of propagation: x-direction.